# Sign changes of Kloosterman sums with moduli having at most two prime factors

### v1

@@ 6.2. Analysis of $N(s)$
$N(s)=\frac{1}{(2\pi i)^{3}}\iiint$ $\check{F}_{\sqrt{D}}(t_{1})\check{F}_{\sqrt{D}}(t_{2})\frac{\sqrt{D}^{t_{1}+t_{2}+w}}{t_{1}+t_{2}+w}$
×$\times\prod_{p\mid\Pi_{l}}\left(1-\frac{1}{p^{t_{1}}}-\frac{1}{p^{t_{2}}}+\frac{1}{p^{t_{1}+t_{2}+w}}\right)$
×$\times\prod_{p\nmid\Pi_{l}}\left(1-\frac{1}{\beta(p,s)p^{t_{1}}}-\frac{1}{\beta(p,s)p^{t_{2}}}+\frac{1}{\beta(p,s)p^{t_{1}+t_{2}+w}}\right)\text{d}t_{1}\text{d}t_{2}\text{d}w.$
6.3. Upper bound for $R_{3}(X)$
From (7) and $N(s)$ we obtained above, there is
::=:= $\frac{X}{(2\pi i)^{4}}\iiiint\tilde{g}(s+1)G(s+1)\zeta^{\kappa}(s+1)\check{F}_{\sqrt{D}}(t_{1})\check{F}_{\sqrt{D}}(t_{2})$
(12) ×$\ \ \ \ \ \ \times H(s,t_{1},t_{2},w)I(s,t_{1},t_{2},w)\frac{X^{s}\sqrt{D}^{t_{1}+t_{2}+w}}{t_{1}t_{2}w}\text{d}s\text{d}t_{1}\text{d}t_{2}\text{d}w,$
where each integral is over $1+it$, $t\in\mathbb{R}$ and
@@ 6.3. Upper bound for $R_{3}(X)$
$I(s,t_{1},t_{2},w):=$ $\prod_{p\mid\Pi_{l}}\left(1-\frac{1}{\beta(p,s)p^{t_{1}}}-\frac{1}{\beta(p,s)p^{t_{2}}}+\frac{1}{\beta(p,s)p^{t_{1}+t_{2}+w}}\right)^{-1}$
×$\times\prod_{p\mid\Pi_{l}}\left(1-\frac{1}{p^{t_{1}}}-\frac{1}{p^{t_{2}}}+\frac{1}{p^{t_{1}+t_{2}+w}}\right)$
@@
Let
$K(s,t_{1},t_{2},w):=$ $H(s,t_{1},t_{2},w)$
×$\times\prod_{p}\left(1-\frac{1}{p^{s+t_{1}+1}}\right)^{-\kappa}\prod_{p}\left(1-\frac{1}{p^{s+t_{2}+1}}\right)^{-\kappa}\prod_{p}\left(1-\frac{1}{p^{s+t_{1}+t_{2}+w+1}}\right)^{\kappa}$
(13) ×$\times\left(\frac{\zeta(s+t_{1}+t_{2}+w+1)}{\zeta(s+t_{1}+1)\zeta(s+t_{2}+1)}\right)^{\kappa}\left(\frac{s(s+t_{1}+t_{2}+w)}{(s+t_{1})(s+t_{2})}\right)^{\kappa},$
@@
$=$ $\frac{X}{(2\pi i)^{4}}\iiiint\tilde{g}(s+1)G(s+1)\zeta^{\kappa}(s+1)\check{F}_{\sqrt{D}}(t_{1})\check{F}_{\sqrt{D}}(t_{2})I(s,t_{1},t_{2},w)$
(14) ×$\ \ \ \ \ \ \times K(s,t_{1},t_{2},w)\left(\frac{(s+t_{1})(s+t_{2})}{s(s+t_{1}+t_{2}+w)}\right)^{\kappa}\frac{X^{s}\sqrt{D}^{t_{1}+t_{2}+w}}{t_{1}t_{2}w}\text{d}s\text{d}t_{1}\text{d}t_{2}\text{d}w.$
We will evaluate the multiple-integral by shifting contours. It can be checked that the quadruple integral appeared in (6.3) and Section 5 in [18] (Page 1220) are the same, except for a finite term $I(s,t_{1},t_{2},w)$. Therefore, after shifting the $w-$contour and $s-$contour through the same way, we only need to handle the integrand with $t_{1}$, $t_{2}$ as
@@
×$\times\check{F}_{\sqrt{D}}(t_{1})\check{F}_{\sqrt{D}}(t_{2})K(0,t_{1},t_{2},0)I(0,t_{1},t_{2},0)(t_{1}t_{2})^{j-1}\frac{\sqrt{D}^{t_{1}+t_{2}}}{(t_{1}+t_{2})^{i+j}}.$
Note that
@@
and $I(0,t_{1},t_{2},0)$ satisfies the condition in Lemma 3.4. Then from Lemma 3.4 with $m=2l$ we have
$=$ $(1+o(1))\frac{4^{2l+1}\tilde{g}(1)I^{(2l)}(0,0,0,0)}{(2l)!}\frac{X}{(\log X)^{2l+1}}\sum_{j=1}^{\kappa}\binom{\kappa}{j}\frac{1}{\Gamma^{2}(j)}$
×$\times\int_{0}^{1}F^{(j+l)}(x)^{2}(1-x)^{j-1}\left(\sum_{i=0}^{j-1}\binom{j-1}{i}(x-1)^{i}4^{j-1-i}\right)\text{d}x$
$=$ $(1+o(1))\frac{4^{2l+1}\tilde{g}(1)I^{(2l)}(0,0,0,0)}{(2l)!}\frac{X}{(\log X)^{2l+1}}$
×$\times\sum_{j=1}^{\kappa}\binom{\kappa}{j}\frac{1}{\Gamma^{2}(j)}\int_{0}^{1}F^{(j+l)}(x)^{2}(1-x)^{j-1}(x+3)^{j-1}\text{d}x.$
In view of the restrictions in (1) again, with the help of mathematical software, we find
$R_{3}(X)\leq(1+o(1))4^{21}2^{10}\left(\prod_{p\mid\Pi_{10}}\log p\right)^{2}\times 8817.853\tilde{g}(1)\frac{X}{(\log X)^{2l+1}},$
@@ Acknowledgements
This work is supported by the National Natural Science Foundation of China (Nos. 12271320, 11871317), and the Natural Science Basic Research Plan for Distinguished Young Scholars in Shaanxi Province of China (No. 2021JC-29).

### v2

@@ 5.2. Analysis of $N(s)$
$N(s)=$ $\frac{1}{(2\pi i)^{3}}\iiint_{(2)}\check{F}_{\sqrt{D}}(t_{1})\check{F}_{\sqrt{D}}(t_{2})\frac{\sqrt{D}^{t_{1}+t_{2}+w}}{t_{1}+t_{2}+w}\frac{1}{\Pi_{l}^{s+1}}\prod_{p\mid\Pi_{l}}\left(1-\frac{1}{p^{t_{1}}}-\frac{1}{p^{t_{2}}}+\frac{1}{p^{t_{1}+t_{2}+w}}\right)$
$\times\prod_{p\mid\Pi_{l}}\left(1+\frac{\kappa}{p^{s+1}}\right)^{-1}\prod_{p\nmid\Pi_{l}}\left(1-\frac{1}{\beta(p,s)p^{t_{1}}}-\frac{1}{\beta(p,s)p^{t_{2}}}+\frac{1}{\beta(p,s)p^{t_{1}+t_{2}+w}}\right)\text{d}t_{1}\text{d}t_{2}\text{d}w.$
5.3. Upper bound for $R_{12}(X)$
From (13) and $N(s)$ we obtained above, we find that
$\sum_{(n,\Pi_{l})=1}g\left(\frac{n\Pi_{l}}{X}\right)\mu^{2}(n)\kappa^{\omega(n)}\left(\sum_{d\mid n\Pi_{l}}\lambda_{d}\right)^{2}$
$:=$ $\frac{X}{(2\pi i)^{4}}\mathop{\iiiint}_{(1)}\tilde{g}(s+1)G(s+1)\zeta^{\kappa}(s+1)\check{F}_{\sqrt{D}}(t_{1})\check{F}_{\sqrt{D}}(t_{2})$
$\ \ \ \ \ \ \times H(s,t_{1},t_{2},w)I(s,t_{1},t_{2},w)\frac{X^{s}\sqrt{D}^{t_{1}+t_{2}+w}}{t_{1}t_{2}w}\text{d}s\text{d}t_{1}\text{d}t_{2}\text{d}w,$
where we shift all contours to $1+it$, $t\in\mathbb{R}$ without passing any poles of the integrand and
@@ 5.3. Upper bound for $R_{12}(X)$
$I(s,t_{1},t_{2},w):=$ $\frac{1}{\Pi_{l}^{s+1}}\prod_{p\mid\Pi_{l}}\left(1-\frac{1}{\beta(p,s)p^{t_{1}}}-\frac{1}{\beta(p,s)p^{t_{2}}}+\frac{1}{\beta(p,s)p^{t_{1}+t_{2}+w}}\right)^{-1}\left(1+\frac{\kappa}{p^{s+1}}\right)^{-1}$
$\times\prod_{p\mid\Pi_{l}}\left(1-\frac{1}{p^{t_{1}}}-\frac{1}{p^{t_{2}}}+\frac{1}{p^{t_{1}+t_{2}+w}}\right)$
@@
Letting
$K(s,t_{1},t_{2},w):=$ $H(s,t_{1},t_{2},w)\left(\frac{\zeta(s+t_{1}+t_{2}+w+1)}{\zeta(s+t_{1}+1)\zeta(s+t_{2}+1)}\right)^{\kappa}\left(\frac{s(s+t_{1}+t_{2}+w)}{(s+t_{1})(s+t_{2})}\right)^{\kappa}$
$\times\prod_{p}\left(1-\frac{1}{p^{s+t_{1}+1}}\right)^{-\kappa}\prod_{p}\left(1-\frac{1}{p^{s+t_{2}+1}}\right)^{-\kappa}\prod_{p}\left(1-\frac{1}{p^{s+t_{1}+t_{2}+w+1}}\right)^{\kappa}$
@@
$=$ $\frac{X}{(2\pi i)^{4}}\mathop{\iiiint}_{(1)}\tilde{g}(s+1)G(s+1)\zeta^{\kappa}(s+1)\check{F}_{\sqrt{D}}(t_{1})\check{F}_{\sqrt{D}}(t_{2})I(s,t_{1},t_{2},w)$
(17) $\ \ \ \ \ \ \times K(s,t_{1},t_{2},w)\left(\frac{(s+t_{1})(s+t_{2})}{s(s+t_{1}+t_{2}+w)}\right)^{\kappa}\frac{X^{s}\sqrt{D}^{t_{1}+t_{2}+w}}{t_{1}t_{2}w}\text{d}s\text{d}t_{1}\text{d}t_{2}\text{d}w.$
We will evaluate the multiple-integral by shifting contours. It can be checked that the quadruple integral appearing in (5.3) and that in Section 5 of [17] (Page 1220) differ by only a finite term $I(s,t_{1},t_{2},w)$. Therefore, after shifting the $w$-contour and $s$-contour in the same way and referring to page 1221, line 17 in [17], we only need to handle the integrand with $t_{1}$, $t_{2}$, where this integrand is
@@
$\times\check{F}_{\sqrt{D}}(t_{1})\check{F}_{\sqrt{D}}(t_{2})K(0,t_{1},t_{2},0)I(0,t_{1},t_{2},0)(t_{1}t_{2})^{j-1}\frac{\sqrt{D}^{t_{1}+t_{2}}}{(t_{1}+t_{2})^{i+j}}.$
Noting that
@@
and $I(0,t_{1},t_{2},0)$ satisfies the condition in Lemma 3.7, we observe that the term
$\prod_{p\mid\Pi_{l}}\left(1-\frac{1}{p^{t_{1}}}-\frac{1}{p^{t_{2}}}+\frac{1}{p^{t_{1}+t_{2}+w}}\right)$
is the essential component in this context. Then from Lemma 3.7 with $m=2l$ we have
$\sum_{(n,\Pi_{l})=1}g\left(\frac{n\Pi_{l}}{X}\right)\mu^{2}(n)\kappa^{\omega(n)}\left(\sum_{d\mid n\Pi_{l}}\lambda_{d}\right)^{2}$
$=$ $(1+o(1))4^{2l+1}\tilde{g}(1)I^{(2l)}(0,0,0,0)\frac{X}{(\log X)^{2l+1}}\sum_{j=1}^{\kappa}\binom{\kappa}{j}\frac{1}{\Gamma^{2}(j)}$
$\times\int_{0}^{1}F^{(j+l)}(x)^{2}(1-x)^{j-1}\left(\sum_{i=0}^{j-1}\binom{j-1}{i}(x-1)^{i}4^{j-1-i}\right)\text{d}x$
$=$ $(1+o(1))4^{2l+1}\tilde{g}(1)I^{(2l)}(0,0,0,0)\frac{X}{(\log X)^{2l+1}}$
$\times\sum_{j=1}^{\kappa}\binom{\kappa}{j}\frac{1}{\Gamma^{2}(j)}\int_{0}^{1}F^{(j+l)}(x)^{2}(1-x)^{j-1}(x+3)^{j-1}\text{d}x.$
Recalling (5), we actually choose $l=1$, $\kappa=2$ in $P(X)$, and this leads to
$I^{(2)}(0,0,0,0)=\frac{(\log p)^{2}}{p}.$
Substituting the above into $R_{12}(X)$ and using
$\sum_{p\leq X}\frac{(\log p)^{2}}{p}=(1+o(1))(\log X)^{2},$
we finally derive
$R_{12}(X)\leq(1+o(1))C_{2}(g,F)\frac{\varepsilon^{2}X}{\log X},$
where
$C_{2}(g,F)=4^{4}\tilde{g}(1)\sum_{j=1}^{2}\binom{2}{j}\frac{1}{\Gamma^{2}(j)}\int_{0}^{1}F^{(j+1)}(x)^{2}(1-x)^{j-1}(x+3)^{j-1}\text{d}x.$
Combining with the lower bound of $R_{11}(X)$ and choosing a sufficiently small $\varepsilon$, we always have
$R_{1}(X)\geq(1+o(1))0.76235\times\widetilde{g}(1)\frac{X}{\log X},$
@@
6. Proof of the Theorem
In view of the restriction (9), we have
$\tau(n;\alpha,\beta)=\alpha_{0}^{3}N(n).$
Applying Proposition 2.1 and Proposition 2.2, we can show that
$\rho R_{1}(X)>2R_{3}(X)$
holds by taking $\rho=16.453\times\alpha_{0}^{3}$. Now if
$\alpha_{0}^{3}N(n)<\rho,$
we further get $\omega(n)\leq 6$ by applying Lemma A.1 below. To establish a sufficiently tight lower bound for $N(7)$, we develop a combinatorial analysis of $N(n)$ in Lemma A.1, where the case $n=7$ requires particular attention due to its non-trivial cardinality constraints. This proves our Theorem.
7. Conclusion
Firstly, we provide the reasons for the selection in Section 2.1. The reason for this selection is that Weil’s bound for Kloosterman sum is $2^{\omega(q)}$ when the modulo $q$ is a square-free number. Therefore, its size is only related to the number of prime factors.
On the one hand, the $\omega(d)=3$” here is difficult to be improved further. Interested readers can try limiting the condition to $\omega(d)=2$”, which cannot improve the conclusion within the framework of this paper. On the other hand, when $\omega(n)=5$, we cannot always select three prime factors so that their product is less than $n^{1/2}$. $1/2$ here comes from the relative size of the variables $Q$ and $MN$ in Lemma 3.3, for which we think this value is also difficult to improve. We will focus on discussing these two points in the end.
In [19], the structure of $\tau\left(n;\alpha,\beta\right)$ is the crux of the whole proof. Compared with [19], our method of selecting $\tau\left(n;\alpha,\beta\right)$ is different. This choice can make the upper bound of $R_{2}(X)$ slightly smaller than before, thereby expanding the allowable range of prime factors in the conclusion. We have provided more detailed explanations in Section 4.
We have proved that Theorem 1.2 is valid when $\omega(q)\leq 6$. For possible improvement of deriving $\omega(q)\leq 5$ under the present framework, we are confronted with two central challenges that require our immediate attention.
Challenge 1: Minimizing $\theta_{1}$ for Prime Factor Triples
Let $\theta_{1}>0$ be a real number satisfying
$p_{i}p_{j}p_{k}\leq q^{\theta_{1}},$
where $p_{i}$, $p_{j}$ and $p_{k}$ are prime factors of $q$. Obviously, $\theta_{1}\leq 1$. Our primary objective is to determine the minimal value of $\theta_{1}$ while ensuring the existence of at least one set of prime factor triples $\left\{p_{i},p_{j},p_{k}\right\}$ that satisfy this inequality. That is, (1) is non-empty. In particular, if Lemma 3.3 remains unchanged, the value of $\theta_{1}$ needs to be 1/2. But for $q$ with only 5 prime factors, such $p_{i}$, $p_{j}$ and $p_{k}$ can not always exist. This minimization process necessitates a meticulous examination of $q$’s prime factorization and the identification of optimal triples that minimize $\theta_{1}$.
However, if we relax the range of $\theta_{1}$ (greater than 1/2) such that the above $p_{i}$, $p_{j}$ and $p_{k}$ can exist, then we still need to solve the following problem, i.e. improving the result of Lemma 3.3.
Challenge 2: Expanding the Range of $\theta_{2}$ Based on Lemma 3.3
Drawing upon the condition in Lemma 3.3, let $\theta_{2}>0$ be a real number satisfying
$Q\leq(MN)^{\theta_{2}}(\log MN)^{-B}.$
From Lemma 3.3 we know that $\theta_{2}=1$. Our second objective is to determine the minimum value of $\theta_{2}$ so that the above inequality still holds.
Appendix A
Lemma A.1.
Let $n>1$ be an integer with $\omega(n)\geq 6$. Then we have
$N(n)_{\omega(n)=6}=10,$
$N(n)_{\omega(n)=7}\geq 20,$
where $N(\cdot)$ is the cardinality of the set $\mathcal{B}(n)$ defined in (1).
Proof.
For $\omega(n)=6$, if $d\mid n$ and $\omega(d)=3$, then necessarily $\omega(n/d)=3$ since $\omega(n)=6$. The number of such divisors $d$ is $\binom{6}{3}=20$. Each unordered pair $\{d,\,n/d\}$ is counted twice, hence $N(n)_{\omega(n)=6}=10$.
Now let $\omega(n)=7$ and write $n=p_{1}p_{2}\cdots p_{7}$ with $p_{1}<p_{2}<\dots<p_{7}$. Because the primes are ordered, the following $14$ triples certainly satisfy $p_{i}p_{j}p_{k}<\sqrt{n}$:
$p_{1}p_{2}p_{3},\;p_{1}p_{2}p_{4},\;p_{1}p_{2}p_{5},\;p_{1}p_{2}p_{6},\;p_{1}p_{3}p_{4},\;p_{1}p_{3}p_{5},\;p_{1}p_{3}p_{6},$
$p_{1}p_{4}p_{5},\;p_{1}p_{4}p_{6},\;p_{2}p_{3}p_{4},\;p_{2}p_{3}p_{5},\;p_{2}p_{3}p_{6},\;p_{2}p_{4}p_{5},\;p_{2}p_{4}p_{6}.$
Denote this set by $\mathcal{G}_{0}$; then $|\mathcal{G}_{0}|=14$.
To improve the lower bound we examine the following six triples
$p_{1}p_{2}p_{7},\;p_{1}p_{3}p_{7},\;p_{1}p_{4}p_{7},\;p_{1}p_{5}p_{6},\;p_{1}p_{5}p_{7},\;p_{1}p_{6}p_{7}.$
If one of them exceeds $\sqrt{n}$, then the complementary product of four primes is at most $\sqrt{n}$, which forces several other triples to be at most $\sqrt{n}$. We proceed by a case analysis, each time assuming that all previous triples from this list are at most $\sqrt{n}$.
Case 1. If $p_{1}p_{2}p_{7}>n^{1/2}$, then $p_{3}p_{4}p_{5}p_{6}\leq n^{1/2}$, whence the six triples
$p_{1}p_{5}p_{6},\;p_{2}p_{5}p_{6},\;p_{3}p_{4}p_{5},\;p_{3}p_{5}p_{6},\;p_{3}p_{4}p_{6},\;p_{4}p_{5}p_{6}$
are all at most $n^{1/2}$. None of these belong to $\mathcal{G}_{0}$, so we obtain $N(n)_{\omega(n)=7}\geq 14+6=20$. If $p_{1}p_{2}p_{7}\leq n^{1/2}$, we keep this triple and proceed.
Case 2. Assume $p_{1}p_{2}p_{7}\leq n^{1/2}$ and $p_{1}p_{3}p_{7}>n^{1/2}$. Then $p_{2}p_{4}p_{5}p_{6}\leq n^{1/2}$, which forces the five triples
$p_{2}p_{5}p_{6},\;p_{3}p_{4}p_{6},\;p_{3}p_{5}p_{6},\;p_{3}p_{4}p_{5},\;p_{4}p_{5}p_{6}\leq n^{1/2}.$
Together with the already counted $p_{1}p_{2}p_{7}$ we have at least $14+1+5=20$ admissible triples; hence $N(n)_{\omega(n)=7}\geq 20$. If $p_{1}p_{3}p_{7}\leq n^{1/2}$, we retain it and continue.
Case 3. Under $p_{1}p_{2}p_{7},p_{1}p_{3}p_{7}\leq n^{1/2}$, suppose $p_{1}p_{4}p_{7}>n^{1/2}$. Then $p_{2}p_{3}p_{5}p_{6}\leq n^{1/2}$, giving the five triples
$p_{1}p_{5}p_{6},\;p_{2}p_{5}p_{6},\;p_{3}p_{4}p_{5},\;p_{3}p_{4}p_{6},\;p_{3}p_{5}p_{6}\leq n^{1/2}.$
Together with the two stored triples we have $14+2+5=21$ admissible triples, so $N(n)_{\omega(n)=7}\geq 21$. If $p_{1}p_{4}p_{7}\leq n^{1/2}$, we keep it.
Case 4. Assume $p_{1}p_{2}p_{7},p_{1}p_{3}p_{7},p_{1}p_{4}p_{7}\leq n^{1/2}$ and $p_{1}p_{5}p_{6}>n^{1/2}$. Then $p_{2}p_{3}p_{4}p_{7}\leq n^{1/2}$, which yields the five triples
$p_{2}p_{3}p_{7},\;p_{2}p_{4}p_{7},\;p_{3}p_{4}p_{5},\;p_{3}p_{4}p_{6},\;p_{3}p_{4}p_{7}\leq n^{1/2}.$
Together with the three stored triples we obtain $14+3+5=22$ admissible triples. If $p_{1}p_{5}p_{6}\leq n^{1/2}$, we add it to the collection.
Case 5. Now suppose $p_{1}p_{2}p_{7},p_{1}p_{3}p_{7},p_{1}p_{4}p_{7},p_{1}p_{5}p_{6}\leq n^{1/2}$ and $p_{1}p_{5}p_{7}>n^{1/2}$. Then $p_{2}p_{3}p_{4}p_{6}\leq n^{1/2}$, forcing
$p_{3}p_{4}p_{5},\;p_{3}p_{4}p_{6}\leq n^{1/2}.$
Together with the four stored triples we have $14+4+2=20$ admissible triples. If $p_{1}p_{5}p_{7}\leq n^{1/2}$, we retain it.
Case 6. Finally, under $p_{1}p_{2}p_{7},p_{1}p_{3}p_{7},p_{1}p_{4}p_{7},p_{1}p_{5}p_{6},p_{1}p_{5}p_{7}\leq n^{1/2}$, assume $p_{1}p_{6}p_{7}>n^{1/2}$. Then $p_{2}p_{3}p_{4}p_{5}\leq n^{1/2}$, whence $p_{3}p_{4}p_{5}\leq n^{1/2}$. This gives $14+5+1=20$ admissible triples.
Contingency. If, after all the previous assumptions, we also have $p_{1}p_{6}p_{7}\leq n^{1/2}$, then all six triples listed above are at most $\sqrt{n}$. None of these lie in $\mathcal{G}_{0}$, so we directly obtain $N(n)_{\omega(n)=7}\geq 14+6=20$.
In every possible case we have shown $N(n)_{\omega(n)=7}\geq 20$. This completes the proof for $\omega(n)=7$.∎
@@ Acknowledgements
This work is supported by the National Natural Science Foundation of China (Nos. 12271320, 11871317), and the Natural Science Basic Research Plan for Distinguished Young Scholars in Shaanxi Province of China (No. 2021JC-29). The authors would like to thank the referees for their very helpful and detailed comments, which have significantly improved the presentation of this paper.